OpenStudy (anonymous):

The perimeter of a rectangle is given by P = 2W + 2L where W is the width and L is the length. The length of a rectangle is 3 more than twice its width. The perimeter of the rectangle is 66 meters. What is the length and width of the rectangle?

OpenStudy (anonymous):

I need help

Vocaloid (vocaloid):

hint: "The length of a rectangle is 3 more than twice its width." means that L = 2W + 3 know where to go from there?

OpenStudy (anonymous):

no still

Vocaloid (vocaloid):

ok, L = 2W + 3 means we can replace "L" with "2W + 3" wherever L shows up

OpenStudy (anonymous):

still confusing

Vocaloid (vocaloid):

so we go from P = 2W + 2L to P = 2W + 2(2W+3) and solve for W

OpenStudy (anonymous):

P= 2W+4W+6 P=6w+6

Vocaloid (vocaloid):

good, and we know that P = 66 so 66 = 6W + 6 what does W equal?

OpenStudy (anonymous):

66=6w+6 -6 -6 60/6=6w/6 10=w

Vocaloid (vocaloid):

good, so that means W = 10, which is the width to find length, we go back to our other equation L = 2W + 3 if W = 10, what is L?

OpenStudy (anonymous):

L=2(10) + 3 L=20+3 L=23

Vocaloid (vocaloid):

good, so width = 10 and length = 23 and that's all